18. Baker Corporation applies manufacturing overhead on the basis of direct labor-hours. At the beginning of...

18.

Baker Corporation applies manufacturing overhead on the basis of direct labor-hours. At the beginning of the most recent year, the company based its predetermined overhead rate on total estimated overhead of $77,250 and 2,500 estimated direct labor-hours. Actual manufacturing overhead for the year amounted to $79,000 and actual direct labor-hours were 2,400.

The overhead for the year was: (Round your intermediate calculations to 2 decimal places.)
Answer #1

Predetermined Overhead rate = Estimated overhead/Estimated total labor hours

= 77,250/2,500 = 30.90 per labor hour

Overhead for the year = 30.90 * 2,400 hours

= 74,160